# Multi-Physics-Enhanced Bayesian Inverse Analysis: Information Gain from Additional Fields

### v1

@@ Abstract
Inverse problems, such as model calibration, are generally ill-posed when treated deterministically. Approaching such problems in a probabilistic, Bayesian sense renders them inherently well-posed and quantifies uncertainty in the solution. This uncertainty quantification is crucial because many complex real-world inverse problems suffer from limited data, often because they rely on measurements of a single physical field. Such data frequently fail to sufficiently reduce parameter uncertainty in Bayesian inverse analysis.
Incorporating easily available data from additional physical fields can substantially decrease this uncertainty. We focus on Bayesian inverse analyses based on computational models, such as those using the finite element method. To incorporate data from additional physical fields, the computational model must be extended to include these fields. While this model extension may have little to no effect on forward model predictions, it can greatly enhance inverse analysis by leveraging the multi-physics data.
Our work proposes this approach of multi-physics-enhanced Bayesian inverse analysis and demonstrates its potential using two models: a scalar-valued model with one-way coupled fields and a complex computational model with fully coupled fields. We quantify the reduction in uncertainty by comparing the effect of single-physics and multi-physics data on the information gain from the prior to the posterior distribution. Our results show that even a few or noisy data points from an additional physical field can considerably increase the information gain, even if the physical field is only weakly or one-way coupled.
Although multi-physics data are often readily available or easily obtainable, it is remarkable that their potential has been largely neglected in (Bayesian) inverse analysis so far. Instead, costly and time-consuming additional experimental setups are often pursued. In contrast, incorporating multi-physics data requires minimal effort when multi-physics models are readily available or straightforward to implement, as is the case with uncoupled and one-way coupled models. This work proposes and promotes the future use of multi-physics-enhanced Bayesian inverse analysis as a cost- and time-saving game-changer across various fields of science and industry, particularly in medicine.
@@ 1 Introduction
Model calibration, which is the identification of model parameters, remains one of the most challenging tasks in all fields of the applied sciences, especially in medicine, and in various industrial applications. The challenges lie in formulating the problem properly, in the need for accurate, efficient, and robust methods, as well as in the availability and quality of the data. The standard approach to identify parameters from experiments and gain quantitative insight is to conduct a specific experiment and use the resulting measurements of obviously relevant quantities. Many excellent examples and powerful methods for this approach can be found in the literature, particularly in the field of computational model calibration, the field on which our work focuses. However, as research problems and computational models have become increasingly complex, this standard approach often faces strict limitations on the type of experiments and measurements that can be conducted at a reasonable cost. This development has led to great difficulties in estimating model parameters. The standard response has been to design new experiments, which are often very expensive and time-consuming. In contrast, a vast number of measurements are readily available, or could be obtained easily and cheaply, but are typically neglected because they are not directly linked to the quantity of interest. In this work, we propose an approach that avoids additional costly and time-consuming experiments and instead leverages data from seemingly unrelated fields.
@@
We illustrate our approach with a simple example. Suppose the task is to take measurements in a certain section of a river to infer some unknown flow boundary conditions. Only very few flow sensors can be placed in the river, which makes it extremely difficult to identify the boundary conditions. Instead of buying additional flow sensors and placing them in hard-to-reach locations, one could simply pour some color into the river. While the color does not affect the flow field, even a few inexpensive, low-resolution images of the color distribution can greatly improve the identification of the boundary conditions. An equivalent approach is standard in gas flows, where smoke is added to a colorless gas to learn more about the flow field. Our approach in this work is similar to this simple idea. We propose not to waste but to use data from additional fields, even if they initially seem irrelevant for the underlying inverse problem. In our simple example above, the (forward) solution of the river flow field from the boundary conditions only asks for solving the Navier-Stokes equations for the flow velocity, which is a single field. However, adding a transported scalar field allows us to also learn from the color distribution, which is highly valuable in the inverse analysis.
@@
Another challenge for the inverse problem of model calibration is the ill-posedness when it is conducted deterministically. This ill-posedness arises due to the fact that every model is merely an approximation of the real system, which the observations capture in a noisy manner. This causes the existence of multiple minima, and even small perturbations in the observations can lead to large changes in the inferred parameters. Moreover, the insight provided by deterministic approaches is limited as they yield only a single value of the parameters. A probabilistic alternative to deterministic inverse analysis is Bayesian inverse analysis. In contrast to the deterministic inverse problem, the Bayesian inverse problem is inherently well-posed in the probabilistic sense. It is also a lot more informative as it provides a full probability distribution over all possible parameters, which quantifies the remaining uncertainty. Most importantly for this work, Bayesian inverse analysis also easily integrates observations from multiple sources or physical fields and can account for different noise levels across these fields.
@@
Note that we use lower-case bold symbols to indicate vectors and lower-case regular symbols to indicate scalars. As model parameters are often hard to measure directly, the inverse analysis of model calibration derives them from measurements of $\bm{y}$. In such cases, Bayesian inverse analysis offers a powerful framework to both include and update a prior belief $p(\bm{x})$ about the parameters $\bm{x}$ based on observations $\bm{y}_{\text{obs}}$ of the model outputs $\bm{y}$. The result of this inverse analysis is the posterior probability distribution $p(\bm{x}|\bm{y}_{\text{obs}})$ according to Bayes’ theorem:
@@
The posterior distribution can remain uninformative, meaning it stays similar to the prior distribution with little change in uncertainty. An uninformative posterior distribution can occur even if we choose an appropriate prior distribution, likelihood model, inverse method, and computational model [1]. In this case, the observations $\bm{y}_{\text{obs}}$ do not sufficiently constrain the range of probable parameters $\bm{x}$. On the one hand, this can result from observations that are noisy or sparse, as depicted in Figure 1. Limited data availability and quality are common challenges across all fields of science and engineering applications. For example, it impedes patient-specific predictions in medical applications. On the other hand, insufficient constraints, and thus an infinite number of solutions for the parameters $\bm{x}$, can also appear irrespective of the number and noise of the observations. For example, the simple model $y=\mathcal{M}(\bm{x}=[x_{1},x_{2}]^{T})=t\cdot(x_{1}+x_{2})$ will never result in a unique solution for $x_{1}$ and $x_{2}$ when observing $y$ over time $t$. This challenge also arises when only integral quantities are observable, as illustrated in Figure 1. For instance, in respiratory medicine, only the integral airflow at the trachea can be easily measured. Local airflow in smaller airways remains inaccessible with current non-invasive clinical methods.
@@
To address these observational challenges, we identify two approaches that reduce uncertainty in an uninformative posterior distribution by enhancing the observations. The first and more common approach is to optimize the experimental design, e.g., by improving the locations of the observations $\bm{y}_{\text{obs}}$. While this approach is the current standard, adjusting real-life experiments is often laborious and costly. In the context of Bayesian optimal experimental design, the optimization further comes with a high computational cost due to the integration over possible designs and repeated evaluation of the posterior distribution [2]. What is more, optimizing the experimental design while observing the same physical field cannot overcome intrinsic observational limitations of this field (compare Figure 1). To overcome these limitations, observations of an additional physical field can provide valuable information and restrict the space of probable input parameters $\bm{x}$. By physical field or simply field, we refer to a physical quantity governed by a distinct set of physical laws. Figure 2 illustrates the difference between this multi-physics-enhanced inverse approach and adjusting the experimental design for a generic inverse analysis. To incorporate the multi-physics observations, the model $\mathcal{M}$ must be extended to model both the original physical field $\bm{y}$ and the additional physical field $\bm{y}_{2}$. It is important to emphasize that this model extension does not aim to improve the forward model predictions. Instead, it seeks to enhance the inverse analysis using the insights from comparing the new model outputs $\bm{y}_{2}$ with the additional observations $\bm{y}_{2,\text{obs}}$. Note that more than one physical field can be added in general. While it may seem trivial that additional observations from coupled physical fields improve a Bayesian inverse analysis, this approach is still rarely applied in practice. During the course of our work, we became aware of one noteworthy exception. This exception is the application of multi-physics-enhanced Bayesian inverse analysis in the domain of geophysics under the term Bayesian joint inversion [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17]. However, in contrast to the context of this work, these Bayesian inverse analyses are not based on coupled but uncoupled fields and do not use models based on discretized differential equations. To the best of our knowledge, only two studies have utilized the potential of multi-physics observations to enhance inverse analysis using a coupled computational model. The first study added displacement observations to electrocardiographic observations to improve the inverse analysis of the electrical state of a heart [18]. This improvement is particularly noteworthy as the second, mechanical field was only weakly one-way coupled to the first, electrical field. The second study probabilistically inferred the local stiffness variation in a ventilated human lung [19]. The authors solved this inverse problem by combining observations of the integral airflow at the ventilator, the first field, with voltage measurements from electrical impedance tomography (EIT), the second, one-way coupled field. Without the additional EIT data containing spatially resolved information, the posterior distribution in this study would have remained uninformative. This brief literature review demonstrates that the use of multi-physics observations to enhance Bayesian inverse analysis of computational models is far from being widely established. We attribute this partially to the implementation costs of multi-physics models, but primarily to a lack of awareness of the potential of this approach. Therefore, the goal of this work is to formally introduce the multi-physics-enhanced approach, to demonstrate its potential, and to encourage researchers and practitioners across various disciplines to adopt it. We particularly propose to leverage the potential of multi-physics-enhanced Bayesian inverse analysis in any of the following scenarios:
@@
In the first two scenarios, the small effort of adopting the multi-physics-enhanced approach can be rewarded with a substantial reduction in posterior uncertainty. While the effort may be higher in the third scenario, the potential reward is even greater if the lack of first-field observations has so far prohibited the inverse analysis. Besides proposing the multi-physics-enhanced approach, our contribution is to raise awareness of its low cost-benefit ratio and to investigate the uncertainty reduction through observations of an additional physical field. Our work is the first to quantify this uncertainty reduction by computing the information gain from the prior to the posterior distribution. In particular, we consider the relative increase in information gain when moving from single-physics to multi-physics-enhanced Bayesian inverse analysis. We analyze this relative increase in information gain for different numbers of observations, noise levels, coupling types, and coupling strengths.
@@ 2 Bayesian inverse analysis with multi-physics observations
We start with an introduction to Bayesian inverse analysis to perform parameter calibration based on computational models. To include multi-physics observations in such analyses, we propose a multi-physics log-likelihood function. In Section 2.1, we provide an overview of coupling types in multi-physics models and their implications for multi-physics-enhanced Bayesian inverse analysis. To avoid unnecessary complexity, our notation does not distinguish between random variables, their realizations, and normal variables.
Model calibration aims to identify model parameters $\bm{x}$ that minimize the discrepancy between the model outputs $\bm{y}=\mathcal{M}(\bm{x})$ and observational data $\bm{y}_{\text{obs}}$. In the context of computational models, the model parameters $\bm{x}$ can represent material parameters, geometric parameters, boundary conditions, and initial conditions. Identifying the parameters $\bm{x}$ from the observations $\bm{y}_{\text{obs}}$ constitutes an inverse problem. From a deterministic perspective, this inverse problem is formulated as an optimization problem, which is often ill-posed. This ill-posedness arises due to the existence of multiple minima or because small perturbations in the observations $\bm{y}_{\text{obs}}$ lead to large changes in the inferred parameters $\bm{x}$. Both effects may result from noise in the observations or from model error, since any model is just an approximation of the real system. Moreover, deterministic solution approaches yield only a single value of the parameters $\bm{x}$. This point estimate may not be a global optimum and neglects alternative solutions without providing an uncertainty estimate. A probabilistic alternative to deterministic inverse analysis is Bayesian inverse analysis. Bayesian inverse analysis reformulates the inverse problem as a statistical inference problem according to Bayes’ theorem (see Equation 2). In contrast to the deterministic inverse problem, the Bayesian inverse problem is inherently well-posed in a probabilistic sense and naturally accounts for observational noise and model error. Its solution is a full probability distribution over all possible parameters $\bm{x}$, which quantifies the remaining uncertainty in the parameters. Most importantly for this work, Bayesian inverse analysis also easily integrates observations $\bm{y}_{\text{obs}}$ from multiple sources or physical fields. We will demonstrate this in the following, after introducing the Bayesian approach in more detail.
@@
During Bayesian inverse analysis, the likelihood distribution $p(\bm{y}_{\text{obs}}|\mathcal{M}(\bm{x}))$ is typically evaluated point-wise for some fixed observations $\bm{y}_{\text{obs}}$. In this case, the likelihood can be seen as a function of the parameters $\bm{x}$. An evaluation of this likelihood function for given parameters $\bm{x}$ involves computing the model output $\bm{y}=\mathcal{M}(\bm{x})$ for this parameter and comparing it to the observations $\bm{y}_{\text{obs}}$. For observations of multiple physical fields, each of the $N_{\text{fields}}$ physical fields is observed at $N_{\text{obs},j}$ spatio-temporal coordinates $\bm{c}_{ij}$. Here, $i$ and $j$ denote the $i$-th observation of the $j$-th field. We assume that each of these observations $\bm{y}_{j,\text{obs},\bm{c}_{ij}}$ is equated with a noisy realization of the corresponding output field $\bm{y}_{j}$ of a multi-physics model $\mathcal{M}(\bm{x})$. We represent a model evaluation of the $j$-th physical field with parameters $\bm{x}$ at the $i$-th observed coordinate $\bm{c}_{ij}$ with $\mathcal{M}_{j}(\bm{x},\bm{c}_{ij})$. To compare model outputs and observations, we need a noise model that describes their statistical relationship. This work assumes independent and identically distributed Gaussian noise with zero mean and a diagonal covariance matrix. We further assume that the field-specific noise $\bm{\varepsilon}_{j}$ is added to the model outputs $\mathcal{M}_{j}(\bm{x}_{\text{true}},\bm{c}_{ij})$ evaluated at the ground-truth model parameters $\bm{x}_{\text{true}}$ to obtain an observation $\bm{y}_{j,\text{obs},\bm{c}_{ij}}$:
@@
The posterior distribution $p(\bm{x}|\bm{y}_{\text{obs}})$ is generally not available in closed form. Instead, only point-wise evaluations of the unnormalized posterior are possible. These evaluations are obtained by computing the product of the prior $p(\bm{x})$ and the likelihood $p(\bm{y}_{\text{obs}}|\mathcal{M}(\bm{x}))$ at specific parameter values $\bm{x}$. In this work, which focuses on the multi-physics enhancement of Bayesian inverse analysis, we approximate the normalized posterior using a simple approach with two steps. First, we evaluate the unnormalized posterior on a grid over the uncertain parameters $\bm{x}$. The grid is chosen to cover the high-density regions of the posterior. This grid choice allows us to approximate the normalization constant (i.e., the denominator of Bayes’ theorem given in Equation 2) in a second step using numerical integration with the trapezoidal rule. We then divide the unnormalized posterior by the normalization constant to obtain an approximation of the posterior. However, the number of model evaluations using this grid-based approach grows exponentially with the number of uncertain parameters. This inference approach is therefore feasible only for inverse problems with few uncertain parameters and a computationally inexpensive model $\mathcal{M}(\bm{x})$. Moreover, the high-density regions of the posterior must be known in advance to choose an appropriate grid.
@@
Instead of grid-based approaches, two main classes of methods are commonly used to solve Bayesian inverse problems. Although a thorough review of these methods is beyond the scope of this work, we provide a very brief summary here with entry points into the literature. Clearly, all of the following methods can be combined with our approach of multi-physics enhancement. The first class of methods for solving Bayesian inverse problems comprises sampling methods. These include Markov chain Monte Carlo (MCMC) methods [20, 21, 22], which aim to generate a sequence of samples from the true posterior distribution. Sampling methods also encompass particle methods [23, 24], which approximate the posterior using a set of weighted samples, called particles in this context. In the sequential case, these methods iteratively propagate particles from the prior to the posterior by repeatedly reweighting, resampling, and moving them. The second class of methods for solving Bayesian inverse problems comprises variational methods [25, 26, 27, 28]. These methods approximate the posterior using a parameterized variational density. The variational parameters of this density are optimized by minimizing a statistical distance measure (typically the Kullback-Leibler divergence) between the variational density and the true posterior. Regardless of the solution method, surrogate models [29, 30, 31, 19] can reduce the computational cost of repeatedly evaluating the model $\mathcal{M}(\bm{x})$ in the likelihood.

### v2

@@ Abstract
Inverse analysis, such as model calibration, often suffers from a lack of informative data in complex real-world scenarios. The standard remedy, designing new experimental setups, is often costly and time-consuming, while readily available but seemingly useless data are ignored.
This work proposes incorporating such data from additional physical fields into the inverse analysis, even when the forward model solves a single-physics problem. A Bayesian framework easily incorporates the additional data and quantifies the resulting uncertainty reduction.
We formally introduce the proposed method, which we denote as multi-physics-enhanced Bayesian inverse analysis. Moreover, this work is the first to quantify the reduction in parameter uncertainty by comparing the information gain from the prior to the posterior when using single-physics versus multi-physics data.
We demonstrate the potential of the proposed method in two exemplary applications. Our results show that even a few or noisy data points from an additional physical field can considerably increase the information gain, even when the physical field is only weakly or one-way coupled.
Overall, this work proposes and promotes the future use of multi-physics-enhanced Bayesian inverse analysis as a cost- and time-saving game-changer across various fields of science and industry, particularly in medicine.
@@ 1 Introduction
Model calibration, which is the identification of model parameters, remains one of the most challenging tasks in all fields of the applied sciences, especially in medicine, and in various industrial applications. The challenges lie in formulating the problem properly, in the need for accurate, efficient, and robust methods, as well as in the availability and quality of the data. The standard approach to identify parameters from experiments and gain quantitative insight is to conduct a specific experiment and use the resulting measurements of obviously relevant quantities. Many excellent examples and powerful methods for this approach can be found in the literature, particularly in the field of computational model calibration, the field on which our work focuses. However, as inverse problems have become increasingly complex, this standard approach often faces strict limitations on the type of experiments and measurements that can be conducted at a reasonable cost. This has added to the already substantial challenges in estimating model parameters. The standard response has been to design new experiments, which are often very expensive and time-consuming. In contrast, a vast number of measurements are often readily available, or could be obtained easily and cheaply. These measurements are typically neglected because they are not directly linked to the quantity of interest. In this work, we propose an approach that avoids additional costly and time-consuming experiments and instead leverages this data from seemingly unrelated fields.
@@
We illustrate our approach with a simple example. Suppose the task is to take measurements in a certain section of a river to infer some unknown flow boundary conditions. Only very few flow sensors can be placed in the river, which makes it extremely difficult to identify the boundary conditions. Instead of buying additional flow sensors and placing them in hard-to-reach locations, one could observe the river’s transport of a substance that does not affect the flow field, e.g., the transport of color. Even a few inexpensive, low-resolution images of the substance or color distribution can greatly improve the identification of the boundary conditions. Our approach in this work is similar to this simple idea. We propose not to waste but to use data from additional fields, even if these fields are irrelevant for the forward simulation task and even if they initially seem irrelevant for the underlying inverse problem. In our simple example above, the (forward) solution of the river flow field from the boundary conditions only asks for solving the Navier-Stokes equations for the flow velocity, which is a single field. However, adding a transported scalar field for the inverse analysis allows us to also learn these boundary conditions from the substance or color distribution.
@@
A particular challenge for the inverse problem of model calibration is the ill-posedness in deterministic formulations. This ill-posedness arises due to several reasons. On the one hand, every model is merely an approximation of the real system, which the observations capture in a noisy manner. As a result, an exact solution generally does not exist for a chosen approximate model and the available measurements. On the other hand, many forward problems governed by partial differential equations have an integral or smoothing character. The corresponding inverse operator then behaves like a differentiation, which strongly amplifies noise in the observations. As a consequence, even small perturbations in the observations can lead to large changes in the inferred parameters. Deterministic inverse methods address this ill-posedness by minimizing a discrepancy measure between model outputs and observations, while adding some form of regularization. Although regularization can help, deterministic inverse problems typically still face the existence of multiple minima and unstable solutions with respect to small changes in the observations. The inverse problem may therefore remain ill-posed with deterministic methods. Moreover, deterministic approaches provide limited insight as they yield only a single value of the parameters and do not quantify the uncertainty of this estimate. A probabilistic alternative to deterministic inverse analysis is Bayesian inverse analysis. In contrast to the deterministic inverse problem, the Bayesian inverse problem is well-posed in the probabilistic sense for most practical applications. By representing the unknown parameters as a random variable, the Bayesian approach yields a probability distribution over all possible parameter values. This probability distribution is a lot more informative than a deterministically derived point estimate and quantifies the remaining uncertainty in the parameters. Most importantly for this work, Bayesian inverse analysis also easily integrates observations from multiple sources or physical fields and can account for different noise levels across these fields.
@@
Note that we use lower-case bold symbols to indicate vectors and lower-case regular symbols to indicate scalars. As model parameters $\bm{x}$ are often hard to measure directly, the inverse analysis of model calibration derives them from measurements of $\bm{y}$. In such cases, Bayesian inverse analysis offers a powerful framework to both include and update a prior belief $p(\bm{x})$ about the parameters $\bm{x}$ based on noisy observations $\bm{y}_{\text{obs}}$ of the model outputs $\bm{y}$. The result of this inverse analysis is the posterior probability distribution $p(\bm{x}|\bm{y}_{\text{obs}})$ according to Bayes’ theorem:
@@
The posterior distribution can remain uninformative, meaning it stays similar to the prior distribution with little change in uncertainty. An uninformative posterior distribution can occur even if we choose an appropriate prior distribution, likelihood model, inverse method, and computational model [undef]. In this case, the observations $\bm{y}_{\text{obs}}$ do not sufficiently constrain the range of probable parameters $\bm{x}$. On the one hand, this can result from observations that are noisy or sparse, as depicted in Figure 1. Limited data availability and quality are common challenges across all fields of science and engineering applications. For example, it impedes patient-specific predictions in medical applications. On the other hand, insufficient constraints, and thus an infinite number of solutions for the parameters $\bm{x}$, can also appear irrespective of the number and noise of the observations. To provide an academic example, the simple model $y=\mathcal{M}(\bm{x}=[x_{1},x_{2}]^{T})=t\cdot(x_{1}+x_{2})$ will never result in a unique solution for the correlated parameters $x_{1}$ and $x_{2}$ when observing $y$ over time $t$. This challenge also arises in practical applications when the model output $\bm{y}$ can only be observed as its integral over a domain, as illustrated in Figure 1. For instance, in respiratory medicine, only the integral airflow at the trachea can be easily measured. Local airflow in smaller airways remains inaccessible with current non-invasive clinical methods.
@@
To address these observational challenges, we identify two approaches that reduce uncertainty in an uninformative posterior distribution by enhancing the observations. The first and more common approach is to optimize the experimental design, e.g., by improving the locations of the observations $\bm{y}_{\text{obs}}$. While this approach is the current standard, adjusting real-life experiments is often laborious and costly. In the context of Bayesian optimal experimental design, the optimization further comes with a high computational cost due to the integration over possible designs and repeated evaluation of the posterior distribution [undefa]. What is more, optimizing the experimental design while observing the same physical field cannot overcome intrinsic observational limitations of this field (compare Figure 1). To overcome these limitations, observations of an additional physical field can provide valuable information and restrict the space of probable input parameters $\bm{x}$. By physical field or simply field, we refer to a physical quantity governed by a distinct set of physical laws. Figure 2 illustrates the difference between this multi-physics-enhanced inverse approach and adjusting the experimental design for a generic inverse analysis. To incorporate the multi-physics observations, the model $\mathcal{M}$ must be extended to model both the original physical field $\bm{y}$ and the additional physical field $\bm{y}_{2}$. It is important to emphasize that this model extension does not aim to improve the forward model predictions. Instead, it seeks to enhance the inverse analysis using the insights from comparing the new model outputs $\bm{y}_{2}$ with the additional observations $\bm{y}_{2,\text{obs}}$. Note that more than one physical field can be added in general.
Upon presentation of this approach, it may seem trivial that incorporating additional observations from coupled physical fields improves a Bayesian inverse analysis. Nevertheless, to our knowledge, this multi-physics-enhanced approach, including a quantification of the information gain, has not yet been directly proposed for the calibration of computational models. We found only the application of a related idea in a different context, namely the approach of Bayesian joint inversion in the domain of geophysics [undefb, undefc, undefd, undefe, undeff, undefg, undefh, undefi, undefj, undefk, undefl, undefm, undefn, undefo, undefp]. However, in contrast to the present work, these Bayesian joint inversions are not based on coupled but uncoupled fields and do not use models based on discretized differential equations. To the best of our knowledge, only two studies have utilized the potential of multi-physics observations to enhance inverse analysis using a coupled computational model. The first study added displacement observations to electrocardiographic observations to improve the inverse analysis of the electrical state of a heart [undefq]. This improvement is particularly noteworthy as the second, mechanical field was only weakly one-way coupled to the first, electrical field. The second study, coming from our own group, probabilistically inferred the local stiffness variation in a ventilated human lung [undefr]. In this study, the inverse problem was solved by combining observations of the integral airflow at the ventilator, the first field, with voltage measurements from electrical impedance tomography (EIT), the second, one-way coupled field. Without the additional EIT data containing spatially resolved information, the posterior distribution in this study would have remained uninformative.
This brief literature review demonstrates that the use of multi-physics observations to enhance Bayesian inverse analysis of computational models is far from being widely established. We attribute this partially to the implementation costs of multi-physics models, but primarily to a lack of awareness of the potential of this approach. Therefore, the goal of this work is to formally introduce the multi-physics-enhanced approach, to demonstrate its potential, and to encourage researchers and practitioners across various disciplines to adopt it. We particularly propose to leverage the potential of multi-physics-enhanced Bayesian inverse analysis in any of the following scenarios, starting from a single-physics inverse analysis:
@@
In the first two scenarios, the small effort of adopting the multi-physics-enhanced approach can be rewarded with a substantial reduction in posterior uncertainty. While the effort may be higher in the third scenario, the potential reward is even greater if the lack of first-field observations has so far prohibited the inverse analysis. Our contribution is to propose the multi-physics-enhanced approach for a broad set of fields and computational models. We further contribute by raising awareness of the low cost-benefit ratio of this approach and by investigating the uncertainty reduction through observations of an additional physical field. Our work is the first to quantify this uncertainty reduction by computing the information gain from the prior to the posterior distribution. In particular, we consider the relative increase in information gain when moving from single-physics to multi-physics-enhanced Bayesian inverse analysis. We analyze this relative increase in information gain for different numbers of observations, noise levels, coupling types, and coupling strengths.
@@ 2 Bayesian inverse analysis with multi-physics observations
We start with an introduction to Bayesian inverse analysis to perform parameter calibration based on computational models. To include multi-physics observations in such an analysis, we propose a multi-physics log-likelihood function. In Section 2.1, we provide an overview of coupling types in multi-physics models and their implications for multi-physics-enhanced Bayesian inverse analysis. To avoid unnecessary complexity, our notation does not distinguish between random variables, their realizations, and normal variables.
Model calibration aims to identify model parameters $\bm{x}$ that minimize the discrepancy between the model outputs $\bm{y}=\mathcal{M}(\bm{x})$ and observational data $\bm{y}_{\text{obs}}$. In the context of computational models, the model parameters $\bm{x}$ can represent material parameters, geometric parameters, boundary conditions, and initial conditions. Identifying the parameters $\bm{x}$ from the observations $\bm{y}_{\text{obs}}$ constitutes an inverse problem. From a deterministic perspective, this inverse problem is formulated as an optimization problem, which is often ill-posed. This ill-posedness arises due to the existence of multiple minima or because small perturbations in the observations $\bm{y}_{\text{obs}}$ lead to large changes in the inferred parameters $\bm{x}$. Both effects may result from noise in the observations or from model error, since any model is just an approximation of the real system. Moreover, deterministic solution approaches yield only a single value of the parameters $\bm{x}$. This point estimate may not be a global optimum and neglects alternative solutions without providing an uncertainty estimate. A probabilistic alternative to deterministic inverse analysis is Bayesian inverse analysis. Bayesian inverse analysis reformulates the inverse problem as a statistical inference problem according to Bayes’ theorem (see Equation 2). In contrast to the deterministic inverse problem, the Bayesian inverse problem is well-posed in a probabilistic sense for most practical applications and naturally accounts for observational noise and model error. Its solution is a full probability distribution over all possible parameters $\bm{x}$, which quantifies the remaining uncertainty in the parameters. Most importantly for this work, Bayesian inverse analysis also easily integrates observations $\bm{y}_{\text{obs}}$ from multiple sources or physical fields. We will demonstrate this in the following, after introducing the Bayesian approach in more detail.
@@
During Bayesian inverse analysis, the likelihood distribution $p(\bm{y}_{\text{obs}}|\mathcal{M}(\bm{x}))$ is typically evaluated point-wise for some fixed observations $\bm{y}_{\text{obs}}$. In this case, the likelihood can be seen as a function of the parameters $\bm{x}$. An evaluation of this likelihood function for given parameters $\bm{x}$ involves computing the model output $\bm{y}=\mathcal{M}(\bm{x})$ for this parameter and comparing it to the observations $\bm{y}_{\text{obs}}$. For observations of multiple physical fields, each of the $N_{\text{fields}}$ physical fields is observed at $N_{\text{obs},j}$ spatio-temporal coordinates $\bm{c}_{ij}$. Here, $i$ and $j$ denote the $i$-th observation of the $j$-th field. We assume that each of these observations $\bm{y}_{j,\text{obs},\bm{c}_{ij}}$ is equated with a noisy realization of the corresponding output field $\bm{y}_{j}$ of a multi-physics model $\mathcal{M}(\bm{x})$. We represent a model evaluation of the $j$-th physical field with parameters $\bm{x}$ at the $i$-th observed coordinate $\bm{c}_{ij}$ with $\mathcal{M}_{j}(\bm{x},\bm{c}_{ij})$. To compare model outputs and observations, we need a noise model that describes their statistical relationship. This work assumes independent and identically distributed (i.i.d.) Gaussian noise with zero mean and a known diagonal covariance matrix. We further assume that the field-specific noise $\bm{\varepsilon}_{j}$ is added to the model outputs $\mathcal{M}_{j}(\bm{x}_{\text{true}},\bm{c}_{ij})$ evaluated at the ground-truth model parameters $\bm{x}_{\text{true}}$ to obtain an observation $\bm{y}_{j,\text{obs},\bm{c}_{ij}}$:
@@
The posterior distribution $p(\bm{x}|\bm{y}_{\text{obs}})$ is generally not available in closed form. Instead, only point-wise evaluations of the unnormalized posterior are possible. These evaluations are obtained by computing the product of the prior $p(\bm{x})$ and the likelihood $p(\bm{y}_{\text{obs}}|\mathcal{M}(\bm{x}))$ at specific parameter values $\bm{x}$. In this work, which focuses on the multi-physics enhancement of Bayesian inverse analysis, we approximate the normalized posterior using a simple approach with two steps. First, we evaluate the unnormalized posterior on a grid over the uncertain parameters $\bm{x}$. The grid is chosen to cover the high-density regions of the posterior. This grid choice allows us to approximate the normalization constant (i.e., the denominator of Bayes’ theorem given in Equation 2) in a second step using numerical integration with the trapezoidal rule. We then divide the unnormalized posterior by the normalization constant to obtain an approximation of the posterior. To integrate a sufficiently smooth function numerically, the trapezoidal rule offers a convergence rate of $\mathcal{O}(1/n^{2})$ in 1D and $\mathcal{O}(1/n)$ in 2D, where $n$ is the number of function evaluations. However, the grid-based approach comes with the drawback that the high-density regions of the posterior must be known in advance to choose an appropriate grid. If the high-density regions are not sufficiently resolved by the grid, the numerical error in the integration will be high. Moreover, the number of model evaluations using this grid-based approach grows exponentially with the number of uncertain parameters. This inference approach is therefore feasible only for inverse problems with few uncertain parameters and a computationally inexpensive model $\mathcal{M}(\bm{x})$.
@@
Instead of grid-based approaches, two main classes of methods are commonly used to solve Bayesian inverse problems. Although a thorough review of these methods is beyond the scope of this work, we provide a very brief summary here with entry points into the literature. Clearly, all of the following methods can be combined with our approach of multi-physics enhancement. The first class of methods for solving Bayesian inverse problems comprises sampling methods. These include Markov chain Monte Carlo (MCMC) methods [undefs, undeft, undefu], which aim to generate a sequence of samples from the true posterior distribution. Sampling methods also encompass particle methods [undefv, undefw], which approximate the posterior using a set of weighted samples, called particles in this context. In the sequential case, these methods iteratively propagate particles from the prior to the posterior by repeatedly reweighting, resampling, and moving them. The second class of methods for solving Bayesian inverse problems comprises variational methods [undefx, undefy, undefz, undefaa]. These methods approximate the posterior using a parameterized variational density. The variational parameters of this density are optimized by minimizing a statistical distance measure (typically the Kullback-Leibler divergence) between the variational density and the true posterior. Regardless of the solution method, surrogate models [undefab, undefac, undefad, undefr] can reduce the computational cost of repeatedly evaluating the model $\mathcal{M}(\bm{x})$ in the likelihood.